OpenStudy (anonymous):

show work for 4x+3y=3;x=32-7y

OpenStudy (agreene):

since you already have a defined x: \(x=32-7y\) it would probably be quickest to plug that in, as such: \[4(32-7y)+3y=3\] solve that to find y, and then use that value to determine x

OpenStudy (anonymous):

substitute x into x 4(32-7y) + 3y= then do distributive property 4 x 32 = 128 and 4 x -7y = -28y 128-28y+3y= then combine like terms -28y + 3y = -25y 128-25y=3 do the opposite orderation 3-128=-125 -25y=-125 do opposite -125/-25 y=5 now substitute y into x equation 32-7(5)=125=x 4(125) + 3(5)=3 500+15=3